OpenStudy (myname):

how do u convert 20 nM to mM

OpenStudy (australopithecus):

is that nm to micro meters?

OpenStudy (australopithecus):

1nm = 0.001um set up a ratio and multiply it by what you are converting make sure the units divide out. \[\frac{0.001um}{1nm}(20nm) = \frac{20nm(0.001um)}{1nm} = \frac{0.001um(20)}{1} = 0.001um(20)\]

OpenStudy (australopithecus):

nm cancels out in the division and you are left with micrometer units, note the u should be considered the greek letter mu or μ